[QUESTION]

39. Suppose you are interested in obtaining a mortgage loan for $250,000 in order to purchase your principal residence. Your lender has suggested that you might be interested in taking an FHA loan. In order to do so, you must pay an additional up-front mortgage insurance

premium (UFMIP) of 1.0% of the mortgage balance. If the interest rate on the fully-amortizing mortgage loan is 5% and the term is 30 years, what is your monthly mortgage payment assuming the UFMIP is financed?

A. $1,342.05

B. $1,355.47

C. $1,498.88

D. $2,500

Ans: B

Difficulty: Intermediate

Learning Objective: 3

[QUESTION]

40. Suppose you are interested in taking an FHA mortgage loan for $350,000 in order to purchase your principal residence. In order to do so, you must pay an additional up-front mortgage insurance premium (UFMIP) of 1.0% of the mortgage balance. If the interest rate on the fully-amortizing mortgage loan is 6% and the term is 30 years and the UFMIP is financed (i.e., it is included in the loan amount), what is the dollar portion of your monthly mortgage payment that is designated to cover the UFMIP?

A. $20.98

B. $291.67

C. $2,119.41

D. $3,500.00

Ans: A

Difficulty: Advanced

Learning Objective: 3
Answer #1

39.

Loan Amount = 250000

UFMIP = 1% = 250000 * 1% = 2500

Total finance = 250000 + 2500 = 252500

Interest rate = 5%

Monthly payments = PXRX (1 + RN [(1 + R)N-1]

P = principal amount = 252500

R = rate of interest = 5% p a or 5%/12 monthly

N = number of installments = 30 years * 12 months = 30 * 12 = 360

Monthly Payments = [252500 * 5%/12 * ((1+5%/12)^360)] / [((1+5%/12)^360)-1]

= [1052.0833 * 4.46774431400616] / [4.46774431400616-1]

= 4700.43933036065 / 3.46774431400616 = 1355.47

Monthly mortgage payment = 1355.47

40.

Loan amount = 350000

UFMIP = 350000 * 1% = 3500

total finance = 350000 + 3500 = 353500

Monthly Interest rate = 6%/12

Number of installments = 30*12 =360

$ amount associated with UFMIP can be calculated in 2 ways

1. calculate monthly payments on total finance of 353500 and deducting monthly payments on mortgage financing of 350000 or

2. calculate monthly payments of UFMIP of 3500

Monthly payments on 3500 = [3500 * 6%/12 * ((1+6%/12)^360)] / [((1+6%/12)^360)-1]

= [17.5 * 6.02257521226289] / 5.02257521226289

= 105.395066214601 / 5.02257521226289

= 20.98

Dollar portion of your monthly mortgage payment that is designated to cover the UFMIP = 20.98

Summary

39. Monthly mortgage payment assuming the UFMIP is financed = B. $1,355.47

40. Dollar portion of your monthly mortgage payment that is designated to cover the UFMIP = A. $20.98